Calculate the net energy change in kilojoules per mole for the formation of KF(s) from the elements: K(s) + 1/2 F2(g) \rightarrow→KF(s). The following information is given: Heat of sublimation for K(s) = 89.2 kJ/mol, Eea for F(g) = –328 kJ/mol Bond dissociation energy for F2(g) = 158 kJ/mol, Ei for K(g) = 418.8 kJ/mol Electrostatic interactions in KF(s) = –821 kJ/mol
Answer #1

Data.

Chemical reaction:

K (s) + 1/2 F2 (g) ==> KF (s); ΔHf = -821 KJ/mol

Heat of sublimation of K (s):

K (s) ==> K (g); ΔHs = +89.2 KJ/mol

Bond dissociation of F2 (g):

1/2 F2 (g) ==> F (g); ΔHd = +158/2 KJ/mol = +79 KJ/mol

Electron affinity of F2 (g):

F (g) + e- ==> F- (g); Eea = -328 KJ/mol

Ionization of K (s):

K (g) ==> K+ (g) + e- ; Ei = 418.8 KJ/mol

-------------------

ΔHnet = -562 KJ/mol